# A Comprehensive Survey on the Trustworthiness of Large Language Models in Healthcare

### v1

@@ 3.6 Explanability
The lack of explainability in LLMs poses a significant barrier to building trust with clinical practitioners, thereby restricting their adoption in real-world healthcare systems. To address this challenge, efforts in clinical practice and research have prioritized improving the transparency of LLMs by developing more effective explanation mechanisms and incorporating human oversight. For instance, Shariatmadari et al. (2024) enhance biomedical applications by integrating knowledge graphs with language models and visualizing attention probabilities to provide clear and interpretable explanations for model predictions. Elsborg and Salvatore (2023) employ local explanation models to generate intuitive, case-specific insights, further advancing the explainability of LLMs and fostering trust in their medical applications.
Another critical area of research focuses on medical imaging explainability. Ghosh et al. (2023) propose a method that iteratively decomposes a black-box (BB) model into interpretable expert models and a residual network, where expert models specialize in specific data subsets and explain their reasoning using First-Order Logic (FOL).
Efforts to improve reasoning and multimodal integration in LLMs have also gained momentum. MedViLaM Xu et al. (2024) and MedThink Gai et al. (2024) focus on improving medical understanding and reasoning by integrating both textual and visual data for complex medical tasks such as question answering and medical image classification, aiming to provide a more holistic view for decision-making. In contrast, MedExQA Kim et al. (2024) prioritizes explainability in medical QA systems by providing multiple explanations for its responses. Causal Graphs Meet Thoughts Luo et al. (2025) and Retrieval and Reasoning on KGs Ji et al. (2024) explore how to enhance complex reasoning in LLMs through knowledge graphs (KGs), helping these models retrieve relevant information from structured sources to answer questions more effectively and with better justification.
For more specialized applications, TOSRR Liu et al. (2025) introduce tree-organized self-reflective retrieval, which aims to improve performance in the niche area of Traditional Chinese Medicine (TCM), combining LLMs with self-reflection techniques for more accurate and contextually relevant responses. Studies like DDCoT Zheng et al. (2023) and Layered Chain-of-Thought Prompting Sanwal (2025) develop methods for multimodal reasoning across multiple agents, with DDCoT focusing on chain-of-thought prompting for structured reasoning across tasks. Lastly, frameworks like A ChatGPT Aided Explainable Framework Liu et al. (2023a) aim for zero-shot diagnosis and interpretation, ensuring more accessible AI support in real-world medical settings.
The trend towards explainability enhancement encompasses both intrinsic and post-hoc approaches. Intrinsic methods, such as knowledge graphs and decomposed expert models, integrate explainability directly into the models, offering natural transparency. Post-hoc methods, including local explanations and textual justifications, elucidate model decisions after predictions, improving user comprehension and trust.
@@ 4 Future Directions
We have reviewed key trust challenges in LLMs and existing solutions. This section highlights current limitations and proposes future directions.
While efforts to enhance LLM truthfulness in healthcare have advanced, gaps remain, including limited adaptability in hallucination mitigation and weak source attribution. Future research should prioritize model-agnostic post-processing, improved self-correction, enhanced uncertainty quantification, and real-time fact-checking via structured knowledge integration.
Existing privacy safeguards, such as de-identification and federated learning, remain imperfect. Strengthening de-identification methods, reinforcing federated learning defenses, refining differential privacy, and exploring homomorphic encryption and real-time audits are crucial next steps.
Safety remains a pressing concern due to adversarial attacks and excessive safety alignment-induced hallucinations. Future work should focus on robust validation mechanisms, refined safety alignment strategies, and comprehensive evaluation frameworks.
Improving robustness remains critical, with adversarial testing and uncertainty quantification methods needing to better handle medical data complexities. Future research should focus on clinically relevant adversarial tests, enhancing uncertainty techniques, and improving instruction robustness.
Despite progress in fairness, key gaps persist, including the need for comprehensive bias assessments and real-world testing of mitigation strategies. Standardizing fairness metrics, conducting real-world evaluations, and assessing long-term impacts on healthcare equity are critical for progress.
Advancements in explainability have yet to bridge significant gaps. Future research should focus on integrating explainability into clinical workflows, developing interactive explanations, and improving multimodal integration to enhance transparency and trust.
Recently, multi-agent frameworks like TriageAgent Lu et al. (2024) have been introduced to streamline complex clinical tasks through agent collaboration. By harnessing their capabilities, we can embed trustworthiness mitigation and evaluation within a multi-agent system, enabling proactive monitoring and intervention across key areas such as truthfulness, privacy, robustness, fairness, and explainability in healthcare LLMs.
Addressing these gaps will ensure that LLMs can be effectively integrated into healthcare systems, improving their reliability, privacy, safety, fairness, and transparency.
@@ 5 Conclusion
The integration of LLMs into healthcare holds great promise, but realizing their full potential requires addressing the critical challenges outlined in this survey. A key concern is truthfulness, as inaccuracies in medical LLMs pose serious risks to patient safety, making their detection and mitigation an ongoing research priority. Equally vital are privacy, safety, robustness, fairness, and explainability, ensuring responsible deployment in real-world clinical settings.
While existing solutions show progress, much work remains to enhance the reliability, transparency, and ethical implications of LLMs in healthcare. Future research must refine these areas, balancing performance with trustworthiness while preventing LLM-based systems systems from worsening healthcare disparities. Comprehensive benchmarks, cross-disciplinary collaboration, and model accountability frameworks will be essential. Additionally, regulatory oversight and ethical guidelines must ensure LLM applications align with medical standards and patient rights.
Ultimately, achieving safe and equitable AI-driven healthcare will require ongoing efforts to improve both technical capabilities and societal frameworks.
@@
Appendix A Comparison of Datasets
We systematically collected and analyzed 41 datasets relevant to the study of trust in LLMs for healthcare. Table 1 provides a comprehensive summary, highlighting key attributes such as data type, content, associated tasks, and the specific trustworthiness dimensions they address. These datasets vary widely, including web-scraped data, curated domain-specific datasets, public text corpora, synthetic data, real-world data, and private datasets. Each dataset’s content specifies its composition, while its associated task defines its primary research application. Additionally, we categorize the datasets based on critical trustworthiness dimensions—truthfulness, privacy and safety, robustness, fairness and bias, and explainability—offering a structured evaluation of their contributions to building reliable and trustworthy healthcare AI.
@@ Appendix A Comparison of Datasets
MultiMedQA Combination of Public and Synthetic Data, Curated Domain-Specific Dataset A benchmark combining six existing medical questions answering datasets spanning professional medicine, research and consumer queries and a new dataset of medical questions searched online, HealthSearchQA. Tasks including Medical Question Answering, Clinical Reasoning, Evidence-Based Medicine, Multilingual and Multimodal Support, Bias and Safety Analysis Fairness and Bias
@@
BiasMedQA Curated Domain-Specific Datasets 1273 USMLE questions Replicate common clinically relevant cognitive biases Fairness and Bias
NEJM Healer Real Data and Curated Domain-Specific Dataset, Consists of Clinical Cases, Diagnostic Pathways, Educational Materials, Interactive Learning Modules Tasks including Diagnostic Skill Development, Medical Education, Simulated Decision-Making, Feedback and Improvement Fairness and Bias
@@
EquityMedQA Curated domain-specific datasets and synthetic data Cover a wide range of medical topics to surface biases that could harm health equity, including implicit and explicit adversarial questions addressing biases like stereotypes, lack of structural explanations, and withholding information. Evaluate the performance of LLMs in generating unbiased, equitable medical responses. Fairness and Bias
”bias” medical dataset curated, domain-specific dataset comprising real-world healthcare data from a large U.S. health system Includes patient demographics, medical histories, and healthcare utilization records. Evaluate and identify racial biases in algorithms used for healthcare management. Fairness and Bias
@@
SQuAD Curated Domain-Specific Dataset Consists of over 100,000 question-answer pairs derived from more than 500 articles from Wikipedia. Each question is paired with a segment of text from the corresponding article, serving as the answer. To develop models that can read a passage and answer questions about it, assessing the model’s ability to understand and extract information from the text. Fairness and Bias
MIMIC Real Data and Curated Domain-Specific Datasets Consists of electronic health records include patient Demographics, Clinical Data, Medical Notes, Treatment Records, Time-series Data Various medical and machine-learning tasks, including Clinical Decision Support, Disease Modeling, Natural Language Processing, Time-series Analysis, Education and Research Fairness and Bias, Explainability
MedQA Curated Domain-Specific Datasets A benchmark that includes questions drawn from the United States Medical License Exam (USMLE). Exam the physicians to test their ability to make clinical decisions Fairness and Bias, Robustness, Explainability
PMC-Patients Curated dataset derived from public text corpora. Contains 167,000 patient summaries extracted from 141,000 PMC articles Designed to benchmark ReCDS systems through two primary tasks: Patient-to-Article Retrieval (PAR), Patient-to-Patient Retrieval (PPR) Robustness
@@
MIMIC- III Public text corpora, real-world data De-identified health-related data from over 40,000 critical care patients, including demographics, vital signs, laboratory tests, medications, and caregiver notes. Epidemiological studies, clinical decision-rule improvement, machine learning in healthcare. Robustness
@@
MedSafetyBench Curated domain-specific dataset and synthetic (generated using GPT-4, Llama-2-7b-chat, and adversarial techniques). 1,800 harmful medical requests violating medical ethics, along with 900 corresponding safe responses. The dataset is structured based on the Principles of Medical Ethics from the American Medical Association (AMA). Assess the medical safety of LLMs by testing whether they refuse to comply with harmful medical requests. Fine-tune LLMs using medical safety demonstrations to enhance their alignment with ethical medical guidelines. Safety
@@
UNIWIZ Synthetic and curated data, including: 17,638 quality-controlled conversations, and 10,000 augmented preference data Features conversations that integrate safety and knowledge alignment. A "safety-priming" method was employed to generate synthetic safety data, and factual information was injected into conversations by retrieving content from curated sources. Fine-tune large language models to enhance their performance in generating safe and knowledge-grounded conversations. Safety
@@
SciFact Curated Domain-Specific Dataset. Includes claims and corresponding evidence abstracts, each annotated with labels indicating whether the claim is supported or refuted, along with rationales justifying the decision. To verify the veracity of scientific claims by identifying supporting or refuting evidence within abstracts and providing justifications for these decisions. Truthfulness
@@
PubHealthTab Curated Domain-Specific Dataset Contains 1,942 real-world public health claims, each paired with evidence tables extracted from over 300 websites. Facilitates evidence-based fact-checking by providing claims and corresponding evidence tables for verification. Truthfulness
@@
LAMA Curated Domain-Specific Dataset. Comprises a set of knowledge sources, each containing a collection of facts. To probe pretrained language models to determine the extent of their factual and commonsense knowledge. Truthfulness
@@
TriviaQA Curated Domain-Specific Dataset. Consists of over 650,000 question-answer pairs, each linked to a set of supporting documents. The questions are sourced from trivia websites, and the answers are derived from the corresponding documents. Training and evaluating models on reading comprehension, specifically focusing on the ability to extract and reason over information from provided documents to answer questions. Truthfulness
@@
Natural Questions (NQ) Real data consists of real anonymized queries from Google’s search engine users, paired with answers derived from entire Wikipedia articles. To develop and evaluate question-answering systems that can read and comprehend entire Wikipedia articles to find answers to user queries. Truthfulness
@@
PopQA Curated Domain-Specific Dataset. consists of 14,000 QA pairs, each associated with fine-grained Wikidata entity IDs, Wikipedia page views, and relationship type information. Designed for open-domain question answering tasks, focusing on evaluating the effectiveness of language models in retrieving and utilizing factual knowledge. Truthfulness
@@
FEVER Curated Domain-Specific Dataset. comprises 185,000 claims, each paired with evidence from Wikipedia articles. These claims are categorized as supported, refuted, or not verifiable. Fact extraction and verification, where models are trained to determine the veracity of claims based on provided evidence. Truthfulness
HEALTHVER Curated Domain-Specific Dataset. Contains 14,330 evidence-claim pairs (SUPPORTS, REFUTES, NEUTRAL) on health claims, mainly COVID-19, verified against scientific articles. Used to train and evaluate models in verifying health claims by classifying them based on scientific evidence. Truthfulness
@@
Med-HALT Synthetic and Real Data, Curated Domain-Specific Dataset, and Public Dataset Consist of Reasoning-Based Assessments, Memory-Based Assessments, Medical Scenarios, Evaluation Metrics Tasks including Evaluation of Hallucination in Medical AI, Reliability Benchmarking, Error Analysis, Mitigation Development Truthfulness
@@
MedICaT Public Text Corpora And Real Data (curated from publicly available biomedical literature) Contains medical images (e.g., radiographs, charts, and diagrams) paired with captions extracted from biomedical literature. Also, includes metadata about the source and context of the images. Task including Medical Image Captioning, Text-Image Retrieval, Medical Reasoning Truthfulness
Med-HallMark Curated Domain-Specific Dataset, , Synthetic and Real Data (includes a mix of real-world medical data and synthetically generated hallucination scenarios) Diverse medical multimodal data, including text, images, and paired annotations. Hierarchically categorized hallucination data, addressing both structural (e.g., object-level) and contextual (e.g., domain knowledge) hallucinations. Task including Hallucination Detection, Hallucination Evaluation, Mitigation Analysis Truthfulness
@@
BioASQ Curated Domain-Specific Dataset; Real Data. The dataset comprises English-language biomedical questions, each accompanied by reference answers and related materials. These questions are designed to reflect real information needs of biomedical experts, making the dataset both realistic and challenging. The primary task is Biomedical Question Answering (QA), which involves systems providing accurate answers to questions based on biomedical data. The dataset supports various QA tasks, including yes/no, factoid, list, and summary questions. Truthfulness
@@
FactualBio Synthetic Data; Public Text Corpora. collection of biographies of individuals notable enough to have Wikipedia pages but lacking extensive detailed coverage. The dataset was generated using GPT-4 and includes biographies of 21 individuals randomly sampled from the WikiBio dataset. Evaluating the factual accuracy of language models, particularly in the context of biography generation. It serves as a benchmark for detecting hallucinations and assessing the factual consistency of generated text. Truthfulness
@@
PubMedQA Curated Domain-Specific Dataset. Consists of over 1,000 question-answer pairs derived from PubMed abstracts, focusing on various biomedical topics. Evaluates the ability of models to comprehend and extract information from biomedical texts to answer specific questions. Truthfulness
@@
MedQuAD Curated Domain-Specific Dataset. The dataset encompasses 37 question types, such as Treatment, Diagnosis, and Side Effects, associated with diseases, drugs, and other medical entities like tests. Designed for medical question answering, the dataset aids in developing and evaluating systems that can understand and respond to medical inquiries. Truthfulness
@@
LiveMedQA2017 Curated Domain-Specific Dataset Consists of 634 question-answer pairs corresponding to National Library of Medicine (NLM) questions Medical question answering, focusing on consumer health questions received by the U.S. National Library of Medicine. Truthfulness
MASH-QA Curated Domain-Specific Dataset. Approximately 25,000 question-answer pairs sourced from WebMD, covering a wide range of healthcare topics. Designed for multiple-answer span extraction in healthcare question answering. Truthfulness
SecureSQL Curated domain-specific dataset Comprises meticulously annotated samples, including both positive and negative instances. The dataset encompasses 57 databases across 34 diverse domains, each associated with specific security conditions. Evaluate and analyze data leakage risks in LLMs, particularly concerning SQL query generation and execution. Privacy
@@
Medical Meadow curated domain-specific dataset It comprises approximately 1.5 million data points across various tasks, including question-answer pairs generated from openly available medical data using models like OpenAI’s Designed to enhance large language models (LLMs) for medical applications Privacy
@@
Electronic Health Records (EHR) at (KHCC) Private dataset gpt-3.5-turbo Clinical research, outcome analysis. Privacy

### v2

@@ 3.6 Explanability
Findings in Explainability Despite advances in rationales and attention maps, most tools lack clinical relevance and faithfulness. Methods often misalign with clinician reasoning; progress needs domain-specific frameworks plus causal/counterfactual explanations.
Lack of explainability limits clinical trust. Recent research explores both intrinsic (model-integrated) and post-hoc (output-interpretation) techniques to make LLM reasoning more interpretable. These methods span a wide range of modalities, including text, graphs, tables, and images, and often incorporate domain-specific knowledge or human-centered reasoning to bridge model outputs and clinical expectations.
Intrinsic explainability methods enhance transparency by aligning model attention with medical knowledge. For example, Shariatmadari et al. (2024) integrate knowledge graphs with attention visualization, while Wang et al. (2022) use Wikipedia-derived medical concepts to guide attention for code prediction, resulting in more concept-consistent outputs. Similarly, structure-to-text models like Pathway2Text Yang et al. (2022) convert biomedical graphs into interpretable narratives, supporting a more intuitive understanding of complex structured inputs.
Post-hoc strategies focus on generating faithful rationales and justifications. FaReBio Fang et al. (2024) highlights how summarization faithfulness suffers with increased abstractiveness and introduces a benchmark to evaluate reasoning fidelity. In the molecular domain, LLM-GCE He et al. (2024) generates counterfactuals for Graph Neural Networks (GNNs) using dynamic feedback to ensure chemically valid, interpretable explanations.
Several methods target zero-shot interpretability without task-specific fine-tuning. RAG-IM Mahbub et al. (2024) enables table-based clinical predictions with natural language justifications, while Liu et al. (2023a) embed ChatGPT into a diagnostic workflow with integrated interpretability components. Retrieval-based systems such as Retrieval + KG Ji et al. (2024) and DDCoT Zheng et al. (2023) further enhance reasoning by chaining knowledge-grounded prompts across modalities.
Explainability in imaging and multimodal contexts is also gaining traction. MedThink Gai et al. (2025) fuses visual and textual inputs to improve multimodal reasoning, and MedExQA Kim et al. (2024) supplies detailed rationales for visual question answering. Ghosh et al. (2023) decomposes black-box decisions into expert modules with first-order logic (FOL) reasoning.
Evaluations
Assessments use attribution heatmaps (e.g., LIME/SHAP), human-in-the-loop ratings, and contrastive/instruction-following tests focused on clarity, factual alignment, and clinical usefulness; standardized healthcare benchmarks remain scarce.
Limitations
Attribution tools have uncertain clinical validity, few studies show improved clinician decisions, and claims are rarely compared across models. Many methods don’t scale to large, multimodal, instruction-following LLMs, leaving faithfulness and practicality unresolved.
3.7 Cost and Efficiency Considerations
While large language models offer transformative potential for healthcare, their real-world deployment faces substantial cost and efficiency constraints. Large medical LLMs (e.g., GPT-4, Med-PaLM 2) are expensive to train, fine-tune, and operate, requiring significant compute, memory, and HIPAA-compliant infrastructure; latency and resource demands further hinder use in low-resource settings. These constraints restrict access to well-funded institutions and slow real-world adoption. Smaller open-source models lower inference costs and enable local/edge deployment but often reduce performance—especially in truthfulness, safety, and robustness. Scaling AI in healthcare therefore requires balancing trustworthiness with computational efficiency.
@@ 4 Future Directions
While core trust dimensions, truthfulness, privacy, robustness, fairness, explainability, and safety, have been the focus of recent work, emerging model paradigms such as multi-agent systems, multi-modal models, and small open-source LLMs introduce new trust challenges underexplored.
Multi-Agent LLMs
Multi-agent LLMs enable distributed reasoning through collaboration between specialized agents, offering improved robustness and self-correction. However, they also raise concerns around coordination, error propagation, and the interpretability of inter-agent communication. Trustworthy multi-agent systems will require protocols for communication, verification, and evaluation that ensure factual alignment and fairness. For example, Lu et al. (2024) introduce TriageAgent, a clinical multi-agent framework with role-specific LLMs for diagnosis and decision-making. While it shows benefits like structured collaboration and early stopping, it also reveals trust challenges, including inconsistent agent confidence, limited transparency, and error propagation—highlighting the need for stronger verification and alignment in high-stakes settings.
Multimodal Foundation Models
Multi-modal LLMs combine text, images, and structured data, better reflecting real-world clinical inputs but complicating trust evaluation. Challenges include cross-modal hallucination, misalignment, and reduced explainability. Addressing these issues will require modality-specific assessments, interpretable fusion strategies, and fairness testing across both textual and visual modalities. For example, Liu et al. (2024b) evaluate open-source multimodal LLMs for genomics and proteomics, highlighting issues with factual consistency and alignment across modalities—underscoring the importance of structured evaluation and interpretable model design in biomedical contexts.
Small Open-Source LLMs
Small open-source medical LLMs are gaining traction for their transparency, adaptability, and lower computational demands, making them attractive for deployment in resource-constrained or privacy-sensitive settings. However, their reduced capacity often leads to increased hallucinations, weaker safety alignment, and heightened privacy risks during fine-tuning on limited clinical data. Ensuring their trustworthiness requires lightweight hallucination mitigation, privacy-preserving training, and scalable evaluation pipelines. Despite their growing use, few studies directly examine these trust issues in small medical LLMs, as most existing research focuses on larger or general-purpose models, leaving a critical gap in the literature.
@@ 5 Conclusion
As large language models continue to expand their role in healthcare, ensuring their trustworthiness remains a critical challenge. This survey reviewed six core dimensions—truthfulness, privacy, safety, robustness, fairness, and explainability—highlighting key methods, benchmarks, and limitations in current research. While recent advances have laid important groundwork, most existing solutions remain narrowly scoped and lack integration across dimensions, limiting their effectiveness in real-world clinical settings.
Acknowledgment
Our work is sponsored by NSF #2442253, NAIRR Pilot with PSC Neocortex and NCSA Delta, Commonwealth Cyber Initiative, Children’s National Hospital, Fralin Biomedical Research Institute (Virginia Tech), Sanghani Center for AI and Data Analytics (Virginia Tech), Virginia Tech Innovation Campus, and generous gifts from Nivida, Cisco, and the Amazon + Virginia Tech Center for Efficient and Robust Machine Learning.
@@ Acknowledgment
Appendix A Inclusion & Exclusion Criteria Details
We conducted an extensive search to identify peer-reviewed papers that address the trustworthiness of LLMs in the healthcare domain. Our search strategy involved a wide range of keyword combinations related to LLMs and core trust dimensions, including: trustworthiness, truthfulness, privacy, safety, robustness, fairness, bias, and explainability. We targeted both domain-specific and general AI venues, focusing on recent publications from 2022 onward.
Specifically, we searched across top-tier conferences and journals, including ACL, EMNLP, NAACL, ICML, NeurIPS, ICLR, KDD, AAAI, IJCAI, Nature, and Science, using platforms such as Google Scholar, Nature, and Science. A full list of keyword queries used in our search is provided below. These queries combined domain terms (medical, clinical) with trust-related dimensions, applied across both “large language models” and “foundation models.” Examples include:
  • large language models, medical, explainability
  • large language models, medical, explainable
  • foundation model, medical, explainability
  • large language models, clinical, explainability
  • large language models, medical, truthfulness
  • large language models, medical, trustworthiness
  • foundation model, medical, trustworthiness
  • large language models, clinical, truthfulness
  • large language models, clinical, safety
  • large language models, medical, safety
  • foundation model, medical, safety
  • large language models, clinical, fairness
  • large language models, medical, fairness
  • foundation model, medical, fairness
  • large language models, clinical, robustness
  • foundation model, medical, robustness
  • large language models, medical, robustness
  • large language models, clinical, privacy
  • large language models, medical, privacy
  • foundation model, medical, privacy
  • large language models, clinical, ethics
  • large language models, medical, ethics
  • foundation model, medical, ethics
In total, our initial search returned approximately 15,322 results, including duplicates and non-relevant papers. Our filtering process proceeded in three stages:
  • Duplicate removal – approximately 11,172 papers eliminated.
  • Relevance screening – we excluded papers that: (a) did not focus on trustworthiness aspects (e.g., architecture design or multi-modal fusion techniques), (b) were not specific to the healthcare domain, or (c) were unpublished preprints (e.g., arXiv manuscripts).
  • Final selection – we curated a final set of 62 papers that directly addressed trust-related challenges in healthcare LLMs, focusing on one or more of the following dimensions: truthfulness, privacy, safety, robustness, fairness, bias, and explainability.
Appendix B Comparison of Datasets
We systematically collected and analyzed 38 datasets relevant to the study of trust in LLMs for healthcare. Table 1 provides a comprehensive summary, highlighting key attributes such as data type, content, associated tasks, and the specific trustworthiness dimensions they address. These datasets vary widely, including web-scraped data, curated domain-specific datasets, public text corpora, synthetic data, real-world data, and private datasets. Each dataset’s content specifies its composition, while its associated task defines its primary research application. Additionally, we categorize the datasets based on critical trustworthiness dimensions—truthfulness, privacy and safety, robustness, fairness and bias, and explainability—offering a structured evaluation of their contributions to building reliable and trustworthy healthcare AI.
@@ Appendix B Comparison of Datasets
MultiMedQA Combination of Public and Synthetic Data, Curated Domain-Specific Dataset 208,000 entries. A benchmark combining six existing medical questions answering datasets spanning professional medicine, research and consumer queries and a new dataset of medical questions searched online, HealthSearchQA. (Med-QA) Tasks including Medical Question Answering, Clinical Reasoning, Evidence-Based Medicine, Multilingual and Multimodal Support, Bias and Safety Analysis Fairness and Bias
@@
BiasMedQA Curated Domain-Specific Datasets 1273 USMLE questions (Med-QA) Replicate common clinically relevant cognitive biases Fairness and Bias
@@
EquityMedQA Curated domain-specific datasets and synthetic data 4,619 examples. Cover a wide range of medical topics to surface biases that could harm health equity, including implicit and explicit adversarial questions addressing biases like stereotypes, lack of structural explanations, and withholding information. (Med-QA) Evaluate the performance of LLMs in generating unbiased, equitable medical responses. Fairness and Bias
@@
SQuAD Curated Domain-Specific Dataset Consists of over 100,000 question-answer pairs derived from more than 500 articles from Wikipedia. Each question is paired with a segment of text from the corresponding article, serving as the answer. (Med-QA)To develop models that can read a passage and answer questions about it, assessing the model’s ability to understand and extract information from the text. Fairness and Bias
@@
MIMIC- III Public text corpora, real-world data De-identified health-related data from over 40,000 critical care patients, including demographics, vital signs, laboratory tests, medications, and caregiver notes. (Med-IE) Epidemiological studies, clinical decision-rule improvement, machine learning in healthcare. Fairness and Bias, Explainability, Robustness
MedQA Curated Domain-Specific Datasets 194,000 multiple-choice medical exam questions. A benchmark that includes questions drawn from the United States Medical License Exam (USMLE). (Med-QA) Exam the physicians to test their ability to make clinical decisions Fairness and Bias, Robustness, Explainability, Truthfulness, Privacy
PMC-Patients Curated dataset derived from public text corpora. Contains 167,000 patient summaries extracted from 141,000 PMC articles (Med-IE) Designed to benchmark ReCDS systems through two primary tasks: Patient-to-Article Retrieval (PAR), Patient-to-Patient Retrieval (PPR) Robustness
MedSafetyBench Curated domain-specific dataset and synthetic (generated using GPT-4, Llama-2-7b-chat, and adversarial techniques). 1,800 harmful medical requests violating medical ethics, along with 900 corresponding safe responses. The dataset is structured based on the Principles of Medical Ethics from the American Medical Association (AMA). (Med-Gen) Assess the medical safety of LLMs by testing whether they refuse to comply with harmful medical requests. Fine-tune LLMs using medical safety demonstrations to enhance their alignment with ethical medical guidelines. Safety
@@
UNIWIZ Synthetic and curated data, including: 17,638 quality-controlled conversations, and 10,000 augmented preference data 17,638 conversations and 10,000 augmented preference data. Features conversations that integrate safety and knowledge alignment. A "safety-priming" method was employed to generate synthetic safety data, and factual information was injected into conversations by retrieving content from curated sources. (Med-Gen) Fine-tune large language models to enhance their performance in generating safe and knowledge-grounded conversations. Safety
@@
SciFact Curated Domain-Specific Dataset. 2,011 claims. Includes claims and corresponding evidence abstracts, each annotated with labels indicating whether the claim is supported or refuted, along with rationales justifying the decision. (Med-Gen) To verify the veracity of scientific claims by identifying supporting or refuting evidence within abstracts and providing justifications for these decisions. Truthfulness
@@
PubHealthTab Curated Domain-Specific Dataset Contains 1,942 real-world public health claims, each paired with evidence tables extracted from over 300 websites. (Med-Gen) Facilitates evidence-based fact-checking by providing claims and corresponding evidence tables for verification. Truthfulness
@@
LAMA Curated Domain-Specific Dataset. 24,223 entries of knowledge sources. Comprises a set of knowledge sources, each containing a collection of facts. (Med-Gen) To probe pretrained language models to determine the extent of their factual and commonsense knowledge. Truthfulness
@@
TriviaQA Curated Domain-Specific Dataset. Consists of over 650,000 question-answer pairs, each linked to a set of supporting documents. The questions are sourced from trivia websites, and the answers are derived from the corresponding documents. (Med-QA) Training and evaluating models on reading comprehension, specifically focusing on the ability to extract and reason over information from provided documents to answer questions. Truthfulness
@@
Natural Questions (NQ) Real data  99.80 GB, with downloaded files accounting for 45.07 GB and the generated dataset occupying 54.73 GB. consists of real anonymized queries from Google’s search engine users, paired with answers derived from entire Wikipedia articles. (Med-QA) To develop and evaluate question-answering systems that can read and comprehend entire Wikipedia articles to find answers to user queries. Truthfulness
@@
PopQA Curated Domain-Specific Dataset. consists of 14,000 QA pairs, each associated with fine-grained Wikidata entity IDs, Wikipedia page views, and relationship type information. (Med-QA) Designed for open-domain question answering tasks, focusing on evaluating the effectiveness of language models in retrieving and utilizing factual knowledge. Truthfulness
@@
FEVER Curated Domain-Specific Dataset. comprises 185,000 claims, each paired with evidence from Wikipedia articles. These claims are categorized as supported, refuted, or not verifiable. (Med-Gen) Fact extraction and verification, where models are trained to determine the veracity of claims based on provided evidence. Truthfulness
@@
HEALTHVER Curated Domain-Specific Dataset. contains 14,330 evidence-claim pairs labeled as SUPPORTS, REFUTES, or NEUTRAL, derived from real-world health claims, mainly about COVID-19, verified against scientific articles. (Med-Gen) Training and evaluating models on the task of verifying the truthfulness of health-related claims by assessing their alignment with scientific evidence. This involves classifying claims as supported, refuted, or neutral based on the provided evidence. Truthfulness
@@
Med-HALT Synthetic and Real Data, Curated Domain-Specific Dataset, and Public Dataset 59,254 entries. Consist of Reasoning-Based Assessments, Memory-Based Assessments, Medical Scenarios, Evaluation Metrics (Med-Gen) Tasks including Evaluation of Hallucination in Medical AI, Reliability Benchmarking, Error Analysis, Mitigation Development Truthfulness
@@
MedICaT Public Text Corpora And Real Data (curated from publicly available biomedical literature) 217,060 figures extracted from 131,410 open-access papers. Contains medical images (e.g., radiographs, charts, and diagrams) paired with captions extracted from biomedical literature. Also, includes metadata about the source and context of the images. (Med-Gen) Task including Medical Image Captioning, Text-Image Retrieval, Medical Reasoning Truthfulness
@@
BioASQ Curated Domain-Specific Dataset; Real Data. 3,743 training questions and 500 test questions. The dataset comprises English-language biomedical questions, each accompanied by reference answers and related materials. These questions are designed to reflect real information needs of biomedical experts, making the dataset both realistic and challenging. (Med-QA) The primary task is Biomedical Question Answering (QA), which involves systems providing accurate answers to questions based on biomedical data. The dataset supports various QA tasks, including yes/no, factoid, list, and summary questions. Truthfulness
@@
FactualBio Synthetic Data; Public Text Corpora. collection of biographies of individuals notable enough to have Wikipedia pages but lacking extensive detailed coverage. The dataset was generated using GPT-4 and includes biographies of 21 individuals randomly sampled from the WikiBio dataset. (Med-Gen) Evaluating the factual accuracy of language models, particularly in the context of biography generation. It serves as a benchmark for detecting hallucinations and assessing the factual consistency of generated text. Truthfulness
@@
PubMedQA Curated Domain-Specific Dataset. Consists of over 1,000 question-answer pairs derived from PubMed abstracts, focusing on various biomedical topics. (Med-QA) Evaluates the ability of models to comprehend and extract information from biomedical texts to answer specific questions. Truthfulness
@@
MedQuAD Curated Domain-Specific Dataset. The dataset encompasses 37 question types, such as Treatment, Diagnosis, and Side Effects, associated with diseases, drugs, and other medical entities like tests. (Med-QA) Designed for medical question answering, the dataset aids in developing and evaluating systems that can understand and respond to medical inquiries. Truthfulness
@@
LiveMedQA2017 Curated Domain-Specific Dataset Consists of 634 question-answer pairs corresponding to National Library of Medicine (NLM) questions (Med-QA) Medical question answering, focusing on consumer health questions received by the U.S. National Library of Medicine. Truthfulness
@@
MASH-QA Curated Domain-Specific Dataset. Approximately 25,000 question-answer pairs sourced from WebMD, covering a wide range of healthcare topics. (Med-QA) Designed for multiple-answer span extraction in healthcare question answering. Truthfulness
SecureSQL Curated domain-specific dataset Comprises meticulously annotated samples, including both positive and negative instances. The dataset encompasses 57 databases across 34 diverse domains, each associated with specific security conditions. (Med-IE) Evaluate and analyze data leakage risks in LLMs, particularly concerning SQL query generation and execution. Privacy
@@
Medical Meadow curated domain-specific dataset It comprises approximately 1.5 million data points across various tasks, including question-answer pairs generated from openly available medical data using models like OpenAI’s (Med-Gen) Designed to enhance large language models (LLMs) for medical applications Privacy
@@
Electronic Health Records (EHR) at (KHCC) Private dataset gpt-3.5-turbo (Med-IE) Clinical research, outcome analysis. Privacy